Based on a study, the average typing speed of women is normally distributed. Karen verified the results by conducting a similar study and found out that the standard error of the sample mean is 9 words per minute using a sample with a variance of 1 215. Determine the variance of the sampling distribution


1
Expert's answer
2022-06-01T11:40:46-0400
"SE=\dfrac{\sigma}{\sqrt{n}}=\dfrac{\sqrt{\sigma^2}}{\sqrt{n}}""n=\dfrac{\sigma^2}{(SE)^2}=\dfrac{1215}{9^2}=15"




"Var(\bar{X})=\dfrac{\sigma^2}{n}=(9)^2=81"